OpenStudy (anonymous):

Consider the following table of the ages of our first six presidents Ages of the first 6 presidents at inauguration Washington 57 J. Adams 61 Jefferson 57 Madison 57 Monroe 58 JQ Adams 57 Find the mean of the data set (57+61+57+57+58+57)/6 = 347/6 = 57.83 Find the standard deviation (showing all work) How many of the presidents’ ages fall within one standard deviation of the mean?

OpenStudy (anonymous):

so to find std dev it is the sq root of the variance and variance = \[1/(n-1) \sum_{i =1}^{6} (x_i - \mu_x)^2 \] and just sq root after

OpenStudy (anonymous):

ok lol so what you do is \[\frac{ (57-57.83)^2+(61-57.83)^2+(57-57.83)^2+(57-57.83)^2+(58-57.83)^2+(57-57.83)^2 }{ 6- 1 }\]

OpenStudy (anonymous):

then square root it and you will get the std dev

OpenStudy (anonymous):

sorry I was off open study, thanks!

OpenStudy (anonymous):

How many of the presidents’ ages fall within one standard deviation of the mean?

OpenStudy (anonymous):

so the standard deviation would be 2.56668

OpenStudy (anonymous):

actually its 1.6020861400062107

OpenStudy (anonymous):

to see if it is 1 std. dev away you should take the numbers, subtract the mean then divide by the std so the first one you do \[\frac{ 57-57.83 }{ 1.602 } = -0.52\] since its greater than -1 and less than 1 might need someone to check on this though